OpenStudy (anonymous):

ps help me..again for this problem the perimiter of rectangle is 228 cm, the lenght is 39 less than twice the with. find the dimension?

OpenStudy (anonymous):

pls. expain..

OpenStudy (anonymous):

perimeter is 228 the l = 2b - 39 so in the formula d 4b - 78 +2b = 228 6b =150 b = 25 and l = 50 - 39 = 11

OpenStudy (anonymous):

the perimeter is the total length of all the sides. in a rectangle u have 2 equal lengths and 2 equal widths

OpenStudy (anonymous):

hey i went wrong somewhere its not possible

OpenStudy (anonymous):

jimmyrep what are your answers ?

OpenStudy (anonymous):

pls explain

OpenStudy (anonymous):

see ann length = 2 * width - 39 sp we know 2*length + 2* width = 228 4*width - 78 + 2*width = 228 6 * width = 150 width = 150/6 = 25 and this way i get length length = 2 *width -39 = 2*25 - 39 = 50 - 39 =11

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

but when i try to get these in perimeter form then perimeter = 2*width + 2*length perimeter = 2*25 + 2*11 perimeter = 50 + 22 = 72 ...I don't know how am i getting this wrong ..there must be some silly mistake

OpenStudy (anonymous):

but when i multiply every side by 4 then i get the answers ... so i would advise you to write length = 44 width = 100

OpenStudy (anonymous):

ok..i wi try it..thank you for your help

OpenStudy (anonymous):

but if you do that then when you put these figure into the length width relation length = 2*width - 39 is not satisfied ... the thing is my method is right but the answers are wrong .. you follow this method and try getting answers ..sorry i know it confuses ..but I can not give you wrong answers too ..sorry ..i was not able to help you all the way ..